to dilute a solution that is 40% saline, a chemist combines it with a solution that is 15% saline. How much of each solution should she use if she needs 150mL of a solution that is 25% saline?
Answer by stanbon(75887)

to dilute a solution that is 40% saline, a chemist combines it with a solution that is 15% saline. How much of each solution should she use if she needs 150mL of a solution that is 25% saline?
--------------------------------
Equation:
salt + salt = salt
0.40x + 0.15(150-x) = 0.25(150)
----
Multiply thru by 100 to get:
40x + 15*150-15x = 25*150
25x = 10*150
x = 60mL (amt. of 40% solution needed)
150-x = 90mL (amt. of 15% solution needed)
==============================================
Cheers,
Stan H.
==========